# Heavy Neutral Leptons without Prejudice

### v1

@@ 1 Introduction
Several mechanisms, such as seesaw models and scotogenic approaches, can generate nonzero neutrino masses [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Beyond neutrino mass, HNLs can also contribute to the generation of the baryon asymmetry of the universe via leptogenesis [20, 21, 22, 23, 24, 25] and may even explain the observed dark matter relic abundance [26, 27, 28]. In particular, introducing three HNLs into a minimal neutrino SM framework could address all these challenges [29, 22].
Instead of focusing on a specific model, this study adopts a model-independent approach to experimentally test HNLs. Many searches have probed high-mass HNLs through their production or prompt decays in colliders [30, 31, 32, 33, 34, 35, 36], see also Refs. [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. In certain parameter regions, HNLs may have substantially displaced decays within detectors, creating a unique displaced vertex signature [59]. Recent studies suggest that these signatures could be targeted in LHC searches with associated charged leptons [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70], via Higgs decays [71, 72, 73, 74, 75, 76, 77], and even in detectors such as LHCb [78]. Similar efforts are underway at DUNE [79, 80], IceCube [81], proposed lepton colliders [82, 83], and SHiP [84, 85]; the latter shows particular promise for HNLs below the $c$-quark mass through meson decays [86]. ATLAS and CMS have already performed some of these analyzes [87, 88, 89, 90, 91].
Additional far detectors at the LHC, such as FASER [92, 93], MoEDAL-MAPP [94, 95], MATHUSLA [96, 97, 98], ANUBIS [99], CODEX-b [100], and FACET [101], have been proposed to further study long-lived particles, with FASER and MoEDAL-MAPP currently operational. The case of HNLs as long-lived particles has been a focus in recent studies [102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112].
@@ 2 Set-up
where $H$ is the SM Higgs doublet, $\widetilde{H}\equiv i\sigma^{2}H^{*}$ (with $\sigma^{2}$ as the second Pauli spin matrix), $L_{\alpha}$ are the SM lepton doublets for $\alpha=e,\mu,\tau$, and $y_{N\alpha}$ are the Yukawa couplings linking HNLs with SM leptons. The mass of each HNL, $m_{N}$, is introduced, but is not explicitly written here, since it depends on whether $\widetilde{N}$ is Dirac or Majorana.
@@
Similar terms can also arise if additional scalar fields, which gain vevs, are present in the model. The physical mass eigenstates of the three SM neutrinos $\nu$ and the $n$ HNLs are then obtained by diagonalizing the full mass matrix in a manner dictated by the model. To satisfy the neutrino oscillation data, at least two HNLs ($n\geq 2$) are necessary [115, 116]. Additionally, fitting sub-eV neutrino masses [117] imposes specific constraints on the model’s parameter space. Diagonalizing to the mass basis introduces mixings between HNLs and SM neutrinos, which, if present, allows HNLs to interact with other SM particles, especially the $W^{\pm}$ and $Z$ bosons. For the type-I seesaw mechanism [10, 11, 12, 13, 14, 15, 16], the mixing angle scales with $y_{N\alpha}\,v_{h}/m_{N}$. However, in inverse seesaw and double-seesaw models [118, 119], an additional mass scale reduces the mixing contributions. In particular, some models can fine-tune mixing contributions to be precisely zero [120, 121, 114]. In contrast, in scotogenic models of neutrino mass, it is the Yukawa coupling to the SM Higgs that vanishes [122, 123]. It is important to note that while some models connect Yukawa couplings and mixing angles, these relationships should not be universally assumed.
To maintain a model-independent approach, we treat Yukawa couplings and mixing angles as separate parameters, being agnostic about their possible interdependence. This approach allows for a broader exploration of the HNL parameter space. Such a set-up can be particularly relevant for displaced decay searches and precision Higgs studies.
@@ 4.2 Displaced Vertices
Figure 3: Sensitivity reach of searches for a displaced decay, for $\sqrt{s}=14$ TeV and $\mathcal{L}=3$ ab-1 are shown by regions inside the contours. The thick line corresponds to the assumption of zero background events, whereas for the thin line 65 events were assumed; see the text for further details. The blue dotted line corresponds to the type-I seesaw relation and the red dotted line corresponds to contour of equal decay width for Higgs and gauge-boson mediated decays. The red bands are in tension with the experimental measurements.

### v2

@@ 1 Introduction
Several mechanisms, such as seesaw models and scotogenic approaches, can generate nonzero neutrino masses [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Beyond neutrino mass, HNLs can also contribute to the generation of the baryon asymmetry of the universe via leptogenesis [20, 21, 22, 23, 24, 25] and may even explain the observed dark matter relic abundance [26, 27, 28, 29]. In particular, introducing three HNLs into a minimal neutrino SM framework could address all these challenges [30, 22].
Instead of focusing on a specific model, this study adopts a model-independent approach to experimentally test HNLs. Many searches have probed high-mass HNLs through their production or prompt decays in colliders [31, 32, 33, 34, 35, 36, 37], see also Refs. [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. In certain parameter regions, HNLs may have substantially displaced decays within detectors, creating a unique displaced vertex signature [60]. Recent studies suggest that these signatures could be targeted in LHC searches with associated charged leptons [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73], via Higgs decays [74, 75, 76, 77, 78, 79, 80, 81], and even in detectors such as LHCb [82]. Similar efforts are underway at DUNE [83, 84], IceCube [85], proposed lepton colliders [86, 87], and SHiP [88, 89]; the latter shows particular promise for HNLs below the $c$-quark mass through meson decays [90]. ATLAS and CMS have already performed some of these analyzes [91, 92, 93, 94, 95].
Additional far detectors at the LHC, such as FASER [96, 97], MoEDAL-MAPP [98, 99], MATHUSLA [100, 101, 102], ANUBIS [103], CODEX-b [104], and FACET [105], have been proposed to further study long-lived particles, with FASER and MoEDAL-MAPP currently operational. The case of HNLs as long-lived particles has been a focus in recent studies [106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118].
@@ 2 Set-up
where $H$ is the SM Higgs doublet, $\widetilde{H}\equiv i\sigma^{2}H^{*}$ (with $\sigma^{2}$ as the second Pauli matrix), $L_{\alpha}$ are the SM lepton doublets for $\alpha=e,\mu,\tau$, and $y_{N\alpha}$ are the Yukawa couplings linking HNLs with SM leptons. The mass of each HNL, $m_{N}$, is introduced, but is not explicitly written here, since it depends on whether $\widetilde{N}$ is Dirac or Majorana.
@@
Similar terms can also arise if additional scalar fields, which gain vevs, are present in the model. The physical mass eigenstates of the three SM neutrinos $\nu$ and the $n$ HNLs are then obtained by diagonalizing the full mass matrix in a manner dictated by the model. To satisfy the neutrino oscillation data, at least two HNLs ($n\geq 2$) are necessary [121, 122]. Additionally, fitting sub-eV neutrino masses [123] imposes specific constraints on the model’s parameter space. Diagonalizing to the mass basis introduces mixings between HNLs and SM neutrinos, which, if present, allows HNLs to interact with other SM particles, especially the $W^{\pm}$ and $Z$ bosons. For the type-I seesaw mechanism [10, 11, 12, 13, 14, 15, 16], the mixing angle scales with $y_{N\alpha}\,v_{h}/m_{N}$. However, in inverse seesaw and double-seesaw models [124, 125], an additional mass scale reduces the mixing contributions. In particular, some models can fine-tune mixing contributions to be precisely zero [126, 127, 120]. In contrast, in scotogenic models of neutrino mass, it is the Yukawa coupling to the SM Higgs that vanishes [128, 129]. It is important to note that while some models connect Yukawa couplings and mixing angles, these relationships should not be universally assumed. This had also been emphasized strongly in our previous study [120].
To maintain a model-independent approach, in the mass basis, we treat the Yukawa couplings, the mixing angles, and the physical HNL masses as separate free parameters, being agnostic about their possible interdependence. This approach allows for a broader exploration of the HNL parameter space. Such a set-up can be particularly relevant for displaced decay searches and precision Higgs studies.
@@ 4.2 Displaced Vertices
Figure 3: Sensitivity reach of searches for a displaced decay, for $\sqrt{s}=14$ TeV and $\mathcal{L}=3$ ab-1 are shown by regions inside the contours. The thick line corresponds to the assumption of zero background events, whereas for the thin line 65 events were assumed; see the text for further details. The blue dotted line corresponds to the type-I seesaw relation. The red bands are in tension with the experimental measurements.